OpenStudy (anonymous):

which function results after applying the sequence of transformations to f(x)=x^5

OpenStudy (anonymous):

reflection over the x axis vertically stretch by a factor of 2 shift down 1 unit a. g(x)=-2x^5 +1 b. g(x)=-2x^5-1 c. g(x)= -2(x-1)^5 d. g(x)= -(2x)^5 -1

OpenStudy (anonymous):

answer is b